OpenStudy (anonymous):

A little help? Easy question! :) Please Explain how opposites are applied when solving one-step equations. How do you know what to use to solve a one-step equation?

OpenStudy (primeralph):

give an example of what you're talking about..........

OpenStudy (anonymous):

like this type of thing: x - 15 = 32

OpenStudy (primeralph):

okay, now what do you mean by 'opposites'?

OpenStudy (anonymous):

like if u need to find the answer in an equation of subtraction u add (the opposite of subtracting)

OpenStudy (anonymous):

if u cant answer this question its fine @ primeralph

OpenStudy (dmezzullo):

Ok so say the equation is 2x+20=30 So since you are adding on the left and you wanna get all the like terms together you would subtract 20 on both sides.

OpenStudy (dmezzullo):

So say there is a subtraction sign u wanna do the opposite to get all the like terms together. does this help any? @GabiMB

OpenStudy (primeralph):

@GabiMB I can answer it, I just don't want to end up confusing you the more. All you simply do is to try to isolate the variable you are trying to solve by moving it's components over.

OpenStudy (anonymous):

lol

OpenStudy (primeralph):

@Luis_Rivera yeah, I felt like an idiot........

OpenStudy (primeralph):

Thing is, the problem is too easy to explain that it seems hard to describe.

OpenStudy (dmezzullo):

It doesn't seem like it's a hard problem i jus wanna kno if i explained it right so i am feelin like an idiot

OpenStudy (anonymous):

Example: x - 12 = 50 Step 1. do the opposite of subtraction which is addition ( With the two numbers you see) Step 2. you got an answer it should be 62 but now lets check your work take 62 and subtract it with 12 what you get will be 50 Step 3. Now you got the answer X = 62

OpenStudy (anonymous):

to sum it up what you guys meant to say is, the opposite of a number is its negative, and if its a rational number( meanin a fraction) is its negative reciprocal

OpenStudy (anonymous):

The answer you are looking for is 47, Because 37 + 15 = 47. ( 47 - 15 = 37)

OpenStudy (anonymous):

:P i put 37 Sorry about that

OpenStudy (anonymous):

32*